OpenStudy (anonymous):

What is x2 − 5x − 7 = 0 plugged into the quadratic formula?

OpenStudy (jim766):

\[\frac{ -b \pm \sqrt{b^2-4(a)(c)} }{ 2a }\]

OpenStudy (anonymous):

I think ax2 + bx + c = 0. ?

OpenStudy (jim766):

A B C 1x^2 - 5x -7 = 0

OpenStudy (anonymous):

yes if given the form \[ax^2+bx+c=0\] then the quadratic equation is \[x= \frac{ -b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (jim766):

\[\frac{ 5\pm \sqrt{(-5)^2-4(1)(-7)} }{ 2 }\]

OpenStudy (anonymous):

Why thank you my good sir

OpenStudy (jim766):

yw, do you understand where it all came from?

OpenStudy (anonymous):

yes I was just having trouble visualizing it